# AGITB: A Signal-Level Benchmark for Evaluating Artificial General Intelligence

### v4

@@ 4 AGITB and ARC: A Comparison
The absence of a unified definition of intelligence has led to AGI benchmarks that often suffer from conceptual ambiguity and limited practical utility. Most treat the AGI system as a black box, offering little insight into the mechanisms underlying generalization. In this regard, AGITB takes a distinct approach. Focusing on the emergence of intelligence from low-level signal processing provides a biologically grounded and mechanistically interpretable framework.
Among existing benchmarks, the Abstraction and Reasoning Corpus (ARC), introduced by Chollet (2019), is perhaps most aligned in spirit, as it also prioritizes generalization over task-specific optimization. ARC presents visual reasoning tasks that require a model to infer novel transformations from a small set of input-output examples—such as recoloring, rearranging, or modifying spatial patterns in a spatial grid.
@@
Task spatial dimensions $2$infty
@@
4.1 Human Performance on ARC
The ARC benchmark claims that its tasks are solvable by humans without specific training, as stated by (Chollet, 2019, p. 46). However, according to the official ARC-AGI Leaderboard [^1] https://arcprize.org/leaderboard, only curated panels of individuals consistently solve all tasks. Individual human performance typically falls short, with reported average accuracy around 77%. Given that ARC tasks are designed to reflect human-like priors, this performance ceiling raises concerns about the benchmark’s reliability as an indicator of AGI. If even humans struggle to achieve perfect performance on a benchmark that presupposes human priors, its relevance for evaluating artificial systems with or without those priors is limited.
@@ 5 Conclusion
AGITB remains solvable by humans yet unsolved by classical algorithms and current state-of-the-art AI systems. This persistent gap provides strong empirical evidence that AGITB captures essential aspects of general intelligence. While the absence of a computational solution does not constitute formal proof of the benchmark’s adequacy, the consistent ability of humans to succeed where machines fail suggests that AGITB effectively distinguishes between narrow and general intelligence. As such, it offers a discriminative test and a principled framework for guiding the development of truly general AI systems.
@@
Appendix A LLM prompt
You are managing two binary-pattern prediction models, ‘A’ and ‘B’, participating in a signal-level AGI test. You will receive a 6-bit binary string, prefixed by a model name (e.g., A 010101, B 111000) that represents a 6-bit sensory input to the corresponding model.
Your task is to design a model that **predicts the next 6-bit input**. Each model has an internal state that only updates after receiving an input. Models’ predictions must always be exactly 6 bits long, and their states can handle hundreds of inputs without losing information. Rules:
  1. 1.
    Upon receiving a 6-bit binary input (e.g., A 110001):
    • Update the corresponding model’s internal state.
    • Make a 6-bit prediction based on the model’s state.
    • Respond with the model’s name and its 6-bit prediction.
  2. 2.
    **State updates only occur with input**.
  3. 3.
    **Models A and B follow the same ground operation principles**.
  4. 4.
    **Correct predictions are critical**:
    • Incorrect predictions will trigger adaptation, updating the model’s internal state to improve future predictions.
    • Models learn and adapt continuously.
  5. 5.
    Response format:
    • Each model’s response is a single line:
      ModelName Prediction
    • Example: A 001010
  6. 6.
    Formatting: Keep responses clean and minimal, without extra explanations or punctuation.
  7. 7.
    Begin by outputting both models’ names and predictions, one per line.
We begin now.

### v5

@@ 4 AGITB and ARC: A Comparison
The absence of a unified definition of intelligence has led to AGI benchmarks that often suffer from conceptual ambiguity and limited practical utility. Most offer little insight into the mechanisms underlying generalization. In this regard, AGITB takes a distinct approach. Focusing on the emergence of intelligence from low-level signal-processing building blocks offers a biologically grounded and mechanistically interpretable framework.
Among competing benchmarks, the Abstraction and Reasoning Corpus (ARC), introduced by Chollet (2019), is perhaps most aligned in spirit, as it also prioritizes generalization over task-specific optimization. ARC presents visual reasoning tasks that require a model to infer novel transformations from a small set of input-output examples—such as recoloring, rearranging, or modifying spatial patterns in a spatial grid.
@@
Task spatial dimensions 222 $\infty$
@@ 5 Conclusion
AGITB is solvable by humans yet unsolved by classical algorithms and current state-of-the-art AI systems. This persistent gap provides strong empirical evidence that AGITB captures essential aspects of general intelligence. While the absence of a computational solution does not constitute formal proof of the benchmark’s adequacy, the ability of humans to succeed where machines fail suggests that AGITB effectively distinguishes between narrow and general intelligence. As such, it offers a discriminative test and a principled framework for guiding the development of truly general AI systems.